the diameter of the of a well is 6ft. and the water is 7ft. how many gallons of water are there in the well, reckoning 7.48 gal. to the cubic foot?
Answer by mananth(12270)

Volume V of cylinder = pi*r^2*h
r= 6/2 =3
h=7
...
V=pi*3^2*h
V= 197.9 ft^3
..
gallons = 197.92/7.48
26.45 gallons